Suppose v1;v2;vm is linearly independent in V and w 2 V. Prove that dim span (v1 +w;v2 +w;:::;vm +w) m1.


1
Expert's answer
2021-06-07T17:26:02-0400

If there are (m+1) linearly independent vectors then dim "V\ge m+1>m ⇒⇐." This is the contrapositive. Let m be dimension of V and Vi E V for I=1 to m+1 be set of linearly independent set.

Hence Vi=I to m are also linearly independent this Vi for I=1 to m is basis for V. We need to prove that. Let let W=span {V1,V2......Vm}, then W"\le" V and dim (W)=dim(V)=m, then W=V.

Since Vm+1 E V, has a unique representation using basis vectors, thus

"V_{m+1}=\displaystyle\sum_{i=1}^m \alpha _i \space V_i"

Hence Vi,i E 1. ...,m+1 are linearly dependent.